OpenStudy (anonymous):

Can (x + 5) / x be simplified?

OpenStudy (shamim):

\[\frac{ x }{ x }+\frac{ 5 }{ x }=1+\frac{ 5 }{ x }\]

OpenStudy (shamim):

is it u r expecting

OpenStudy (anonymous):

I worked out a problem and got the answer of x + 5 / x. The instructions were to simplify. I didn't know if the x's could cancel out for an answer of 5.

OpenStudy (anonymous):

shamim you there?

OpenStudy (shamim):

ya i m here